write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0. The solutions are -10, 7
Answer by rwm(914)

(x+10)*(x-7)=0
x^2+10x-7x-70=0
x^2+3x-70=0